OpenStudy (anonymous):

\[\sqrt[3]{343}\] simplify

OpenStudy (anonymous):

7

OpenStudy (anonymous):

please show the work

OpenStudy (anonymous):

\[\sqrt[4]{256}\]

OpenStudy (anonymous):

and this one too

OpenStudy (anonymous):

\[\sqrt[3]{7^{3}}==7\]

OpenStudy (anonymous):

is there any tips to find the number?

OpenStudy (anonymous):

there are no tips just a mind calculations

OpenStudy (anonymous):

\[\sqrt[4]{4^{4}}==\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

you need to know the squares,cubes and fourth powers of at least the numbers 1-10

OpenStudy (anonymous):

so you just memorize it?

OpenStudy (anonymous):

pretty much, yes.

OpenStudy (anonymous):

i see thx